INFRARED EMISSION SPECTRA OF HOT BeF_{2} AND MgF_{2}

\underline{SHANSHAN YU}, ALIREZA SHAYESTEH, and PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; JACEK KOPUT, Department of Chemistry, Adam Mickiewicz University, Grunwaldzka 6, 60-780 Poznan, Poland.

High resolution infrared emission spectra of hot BeF_{2} in the 800–2500 cm^{−1} region have been rotationally analyzed. The ν_{3} fundamental band, ν_{1} +ν_{2}, ν_{1} +ν_{3} and 2ν_{2} +ν_{3} combination bands, and more than 15 hot bands were assigned. The ν_{1} (σ_{g}), ν_{2} (π_{u}) and ν_{3} (σ_{u}) frequencies were directly obtained by fitting several hot bands and combination bands together. A complete analysis of all the bands is in progress and will be presented. High level ab initio calculations have been performed for the MgF_{2} molecule to predict its vibration-rotation spectrum. An infrared emission spectrum of MgF_{2} was recorded and will be assigned with the aid of ab initio calculations.